# for Web Agents

### v1

@@ Abstract
The proliferation of web agents necessitates advanced navigation and interaction strategies within complex web environments. Current models often struggle with efficient navigation and action execution due to limited visibility and understanding of web structures. Our proposed R2D2 framework addresses these challenges by integrating two paradigms: Remember and Reflect. The Remember paradigm utilizes a replay buffer that aids agents in reconstructing the web environment dynamically, thus enabling the formulation of a detailed “map” of previously visited pages. This helps in reducing navigational errors and optimizing the decision-making process during web interactions. Conversely, the Reflect paradigm allows agents to learn from past mistakes by providing a mechanism for error analysis and strategy refinement, enhancing overall task performance. We evaluate R2D2 using the WebArenabenchmark, demonstrating substantial improvements over existing methods, including a 50% reduction in navigation errors and a threefold increase in task completion rates. Our findings suggest that a combination of memory-enhanced navigation and reflective learning promisingly advances the capabilities of web agents, potentially benefiting various applications such as automated customer service and personal digital assistants.
R2D2: Remembering, Reflecting and Dynamic Decision Making
for Web Agents
Tenghao Huang1thanks: Work done during Tenghao’s internship at IBM research.  Kinjal Basu2  Ibrahim Abdelaziz2 Pavan Kapanipathi2  Jonathan May1  Muhao Chen3 1University of Southern California, 2IBM Research, 3University of California, Daivs tenghaoh@usc.edu
@@ 1 Introduction
Web agents—–autonomous AI agents designed to navigate and perform natural language-described tasks within web environments—-have become increasingly integral to applications such as online customer service Huang et al. (2024a), automated data retrieval Huang et al. (2024b), and personalized digital assistants. [^1] https://www.anthropic.com/news/3-5-models-and-computer-use. These agents interact with complex web interfaces to execute user-described tasks, often emulating human actions like clicking buttons, filling forms, and extracting information Shi et al. (2017); Liu et al. (2018); Yao et al. (2022); Zhou et al. (2023).
Figure 1: Traditional methodologies conceptualize web navigation within the framework of an Unknown MDP. The ReACT agent operates under high uncertainty due to incomplete information regarding the outcomes of their actions, leading to erroneous navigational paths, impeding effective task resolution. R2D2 transforms the task into a Known MDP, improving robustness.
Figure 2: This diagram represents various approaches for web agents framed as a search problem, where each node symbolizes a webpage. (a) Reactive: The agent chooses the best immediate actions without any proactive strategy. (b) Tree-search with reflections: the agent investigates different routes by actively navigating websites and allows for reversing direction (shown by dashed arrows). Both a) and b) approaches are Unknown-MDP-based. At each timestep, agents’ observation space is constrained, which typically results in suboptimal or inefficient results. c) R2D2: Our proposed framework constructs the search space leveraging stored state information from a replay buffer. By transforming the task into a Known MDP, R2D2 enhance agents’ ability to navigate and interact with web interfaces.
Despite recent advancements in web agents’ capabilities, a persistent challenge remains: agents frequently fail to navigate effectively within intricate web environments. As illustrated in Figure 1, the successful resolution of a user query usually requires a long series of actions.
The fundamental challenges associated with previous methodologies are twofold: first, these approaches model web navigation as an Unknown Markov Decision Process (MDP), wherein the agent has limited visibility into the consequences of its actions, often leading to suboptimal performance outcomes. Second, prior methods engage in complex reasoning during the inference phase while observing a stream of experiences, and in their simplest forms, they discard incoming trajectories immediately after a single episode. This rapid forgetting of potentially valuable experiences impedes the agent’s capacity to leverage useful information for future decision-making.
@@
Inspired by cognitive studies showing that humans excel at complex tasks by iteratively refining strategies based on feedback Palenciano et al. (2021); Zenkri et al. (2024), as well as by approaches in robotics for structured exploration of unfamiliar spaces Thrun (2002), we propose the R2D2 (Remembering, Reflecting, and Dynamic Decision Making) framework that enhances both navigation and task execution for web agents. Our method transforms the task from Unknown-MDP into a Known MDP by introducing the Remember Paradigm. It leverages a structured replay buffer of the agent’s experiences that guides the agent to more promising avenues Blundell et al. (2016); Schaul et al. (2016); Parisotto and Salakhutdinov (2017); Savinov et al. (2018). At a high level, our approach enables the agent to record and recall previously visited pages—essentially constructing a dynamic map of the web environment—and then leverage this knowledge to refine its strategies. By converting the agent’s experience into a well-organized search space, we empower it to identify reliable navigation routes to target resources rather than re-deriving them from scratch during inference. This reduces the computational overhead at test time and helps avoid repetitive or unproductive exploration.
@@
Our proposed method diverges from traditional techniques by providing a more comprehensive and structured representation of the agent’s historical experiences. Instead of simply recalling past states Agashe et al. (2024) or relying on on-the-fly reasoning Koh et al. (2024); Zhou et al. (2024), our approach organizes the agent’s accumulated experiences into a coherent and reusable resource that effectively guides future decisions. We evaluate our approach using the WebArena benchmark Zhou et al. (2023), where it achieves substantial gains compared to baseline models, including approximately a 50% reduction in navigation errors and a threefold increase in overall task completion rates. Moreover, R2D2 outperforms state-of-the-art methods by 17%, thereby demonstrating a more robust and informed capability for executing complex web-based tasks.
@@ 2 Related Works
Enhancing Web Agents. Existing research has illuminated that language models, without intervention, struggle to express linguistic intent in formal instruction that can control an extra-linguistic environment, such as web site navigation Shi et al. (2017); Liu et al. (2018); Yao et al. (2022); Zhou et al. (2023); Deng et al. (2023). This inadequacy stems primarily from the intrinsic challenges associated with perception, strategic planning, and task execution in the intricate web environments.
To mitigate these challenges, enhancements to web agents have been categorized into two following principal strategies. (1) Perception Alignment: This strategy aims to augment agents’ capabilities in interpreting graphical user interface elements by integrating multimodal data from webpages, enhancing both textual and visual comprehension Gou et al. (2024); Anonymous (2024). (2) Post-hoc Reflection: Studies indicate that enabling agents to engage in reflective practices post interaction can facilitate learning from historical trajectories, thereby improving future task executions Shinn et al. (2023); Song et al. (2024); Pan et al. (2024); Wang et al. (2024a). (3) Online Search Algorithms: This involves the adoption of sophisticated search algorithms, including Monte Carlo Tree Search and other tree-based exploration methods, integrated with high-level planning driven by LLM-derived knowledge Zhang et al. (2024); Koh et al. (2024); Meng et al. (2024). Furthermore, Gu et al. 2024 discusses speculative planning that leverages simulations of world models.
Despite these enhancements, the performance of current web agents are constrained by the assumptions of Unknown MDP, where the potential outcomes of actions are not available. In contrast, this paper proposes a novel approach where we reconstruct the web environment’s structure based on the agents’ exploratory actions, thereby furnishing them with outcome information crucial for making informed and grounded decisions.
Continual Exploration of the Agentic Environment. Tasking agents to explore an unknown environment has been an active research direction in the community Brohan et al. (2023). Recent studies have focused on how agents abstract experiences into actionable skills, a development that is becoming increasingly central to advancements in this field Wang et al. (2023a, b); Liu et al. (2024). Within the domain of web-based agents, these skills are often conceptualized as workflows. Sodhi et al. (2024) have introduced a novel framework that leverages human-engineered workflows to compose policies to tackle web tasks. Although improving agents’ performance, manually crafting workflows can be a tedious process.
@@
Figure 3: An overview of R2D2 architecture, highlighting the Remember and Reflect Paradigms. The Remember paradigm constructs a structured replay buffer from previous observations, enabling the agent to utilize past episodic data through A* search for navigation. Meanwhile, the Reflect paradigm diagnoses errors and generates corrective insights, which are then stored in a reflective memory for future decision-making processes.
@@ 3 Method
In this section, we present R2D2, a framework for tackling complex web navigation tasks by integrating two paradigms: Remember and Reflect. The Remember paradigm constructs a structured replay buffer from past observations(Section 3.2), while the Reflect paradigm diagnoses and corrects errors in failed trajectories(Section 3.3). We then introduce mechanisms of the reflective memory (Section 3.4). Finally, we illustrate how these paradigms interact to improve the agent’s performance through retrieval and in-context learning demonstrations (Section 3.5).
@@ 3.1 Method Overview
Given a user’s task query $q$ and an initial observation $o_{0}$ from the environment, the agent must produce a sequence of actions to address $q$. We define an episode as the process where the agent starts from $o_{0}$ and executes a trajectory $t$. Let $t=\{a_{1},a_{2},\ldots,a_{H}\}$ be the trajectory of length $H$, where each $a_{h}$ is an action at step $h$. After each action $a_{h}$, the agent receives an observation $o_{h}$, thereby forming an observation sequence $O=\{o_{1},o_{2},\ldots,o_{H}\}$. Consider $N$ distinct user queries $\{q_{1},q_{2},\ldots,q_{N}\}$, each associated with its own episode and and observation sequence $O_{i}=\{o_{i,1},o_{i,2},\ldots,o_{i,H}\}$. Across all $N$ user queries, R2D2 forms the union of these observations:$\mathcal{O}_{\text{all}}=\bigcup_{i=1}^{N}O_{i}.$
R2D2 addresses errors in trajectories by categorizing them into two distinct types: (1) Navigation failure. We define $O^{*}$ as the key observations essential for successfully addressing the query $q$. After performing the trajectory $t$, the observation sequence do not contain all the key observations, which leads to agents’ navigation failure in the environment, formally $O\cap O^{*}\neq O^{*}$. The agent fails because of incomplete information or tools to address the user query. (2) Execution failure. After performing $t$, $O\cap O^{*}=O^{*}$. In other words, navigation was successful, since the proper path was followed, but the agent still failed to address the user query.
The Remember paradigm aims to build a replay buffer from $\mathcal{O}_{all}$, allowing the agent to store and revisit past observations and experiences Blundell et al. (2016); Schaul et al. (2016); Parisotto and Salakhutdinov (2017); Savinov et al. (2018). The Reflection paradigm then corrects trajectories that failed due to execution issues, providing explicit rationales for these execution failures. The successful and corrected trajectories and their rationales are stored in a reflective memory for agent’s future reference. Finally, the Retriever leverages this reflective memory by selecting relevant corrected trajectories as in-context demonstrations, thereby continually improving the agent’s performance.
@@ 3.2 Remember Paradigm
Building the Replay Buffer. To effectively represent the web environment and enable its systematic reconstruction from the observation sequence $\mathcal{O}_{\text{all}}$, R2D2 structures the replay buffer as a directed graph $G=(O,E)$. Here, the vertex set $O$ includes the root node $o_{0}$, which corresponds to the homepage observation, and each subsequent vertex represents another webpage observation $o_{i}$. The edge set $E$ consists of tuples $((o_{i},o_{j}),a)$ where each edge corresponds to an action $a$ that transitions the agent from one observation $o_{i}$ to another observation $o_{j}$. Due to the noisy and dynamic nature of web pages, R2D2 store the differences between consecutive observations at each vertex rather than the full webpage state.
$\mathbf{A}^{*}$ Search within the Buffer. The search algorithm employs a breadth-first search ($A^{*}$) strategy Hart et al. (1968); Meng et al. (2024) to navigate and evaluate web environments effectively. Instead of expanding nodes level-by-level, R2D2 incorporates a heuristic that guides the search toward potentially relevant and promising webpages more efficiently Bonet and Geffner (1999); Guez et al. (2018); Moldovan and Abbeel (2012). This heuristic, provided by the LLM, estimates the relevance and utility of exploring a particular webpage node, thereby reducing unnecessary expansions and focusing on paths that are more likely to yield correct information.
@@
0:  Optimal solution trajectory $P^{*}$ for query $q$
@@ 3.3 Reflect Paradigm
R2D2 first determines [^2] Please refer to Figure 7 in Appendices for the detailed prompt. the error type of a failed trajectory $t$. For all navigation failures, R2D2 performs search within the replay buffer to correct their navigation behaviors. We discuss details of how we address navigation failure in Section 3.2. We now discuss using reflection techniques to address execution errors.
@@ 3.4 Reflective Memory Mechanism
We hereby introduce the reflective memory mechanism that stores corrected and truncated trajectories for future retrieval. The reflective memory is structured as a key-value store:
@@ 3.5 Paradigm Coordination and Inference
Exploration Phase. Using a ReACT agent Yao et al. (2023), R2D2 processes user queries and collects observational data to build $\mathcal{O}_{\text{all}}$. Trajectories are classified and corrected via Remember and Reflect paradigms, then stored in the memory.
@@ 4.1 Experimental Setup
ReACTYao et al. (2023) 13.1%
Tree-Search Koh et al. (2024) 16.5% 10.5% 28.1% 25.8% 13.3% 19.2%
@@
Benchmark. We use the WebArena benchmark Zhou et al. (2023). This benchmark comprises diverse web interaction scenarios, ranging from web shopping to customer relationship management system (CMS). The dataset consists of 812 user queries with annotated ground truth trajectories. The Webarena benchmark further provides a set of validators to programmatically validate the functional correctness of each task.
@@
Baselines. We compare our framework against several represetative agent frameworks: (1) ReACT Yao et al. (2023): a widely-used framework, which takes an observation of the environment as input, performs Chain-of-Thought reasoning, and then generates the next action. (2) Tree-Search Koh et al. (2024): an inference-time tree-search strategy to perform best-first tree-search in web environments. It enables agents to revert to the most recently validated state upon encountering a failed trajectory. (3) LATS Zhou et al. (2024): a method based on Monte Carlo tree search that employs LLMs as agents, value functions, and optimizers for decision-making. (4) Anticipatory Reflection Wang et al. (2024a): a method that explicitly considers potential failures before action, alignment and backtracking after actions to maintain task objectives. (5) AutoEval Pan et al. (2024): methods that boost agent performance using domain-general automatic evaluators. (6) BrowserGym Chezelles et al. (2024): a framework that incorporates additional actions and observation tools for agents to interact with the environment. [^5] We also notice there are other methods that use different setups, such as SteP Sodhi et al. (2024) that employs human-engineered workflows, and AWM Wang et al. (2024b) that uses BrowserGym framework that customizes a larger action space than standard WebArena to interact. For the sake of direct comparison, these frameworks that depend on additional human efforts are not taken into comparison here.

### v3

# with a Reflective Agentic Memory

@@ Abstract
The proliferation of web agents necessitates advanced navigation and interaction strategies within complex web environments. Current models often struggle with efficient navigation and action execution due to limited visibility and understanding of web structures. Our proposed R2D2 framework addresses these challenges by integrating two paradigms: Remember and Reflect. The Remember paradigm uses a replay buffer that aids agents in reconstructing the web environment dynamically, thus enabling the formulation of a detailed “map” of previously visited pages. This helps in reducing navigational errors and optimizing the decision-making process during web interactions. Conversely, the Reflect paradigm allows agents to learn from past mistakes by providing a mechanism for error analysis and strategy refinement, enhancing overall task performance. We evaluate R2D2 using the WebArena benchmark, demonstrating substantial improvements over existing methods, including a 50% reduction in navigation errors and a threefold increase in task completion rates. Our findings suggest that a combination of memory-enhanced navigation and reflective learning promisingly advances the capabilities of web agents, potentially benefiting various applications such as automated customer service and personal digital assistants.
R2D2: Remembering, Replaying and Dynamic Decision Making
with a Reflective Agentic Memory
Tenghao Huang1thanks: Work done during Tenghao’s internship at IBM research.  Kinjal Basu2  Ibrahim Abdelaziz2 Pavan Kapanipathi2  Jonathan May1  Muhao Chen3 1University of Southern California, 2IBM Research, 3University of California, Davis tenghaoh@usc.edu
@@ 1 Introduction
Web agents—autonomous AI agents designed to navigate and perform natural language-described tasks within web environments—have become increasingly integral to applications such as online customer service Huang et al. (2025), automated data retrieval Huang et al. (2024), and personalized digital assistants. [^1] https://www.anthropic.com/news/3-5-models-and-computer-use These agents interact with complex web interfaces to execute user-described tasks, often emulating human actions like clicking buttons, filling forms, and extracting information Shi et al. (2017); Liu et al. (2018); Yao et al. (2022); Zhou et al. (2024b).
Figure 1: Traditional methodologies conceptualize web navigation within the framework of an Unknown MDP. The ReACT agent operates under high uncertainty due to incomplete information regarding the outcomes of its actions, leading to erroneous navigational paths and impeding effective task resolution. R2D2 transforms the task into a Known MDP, improving robustness.
Figure 2: This diagram represents various approaches for web agents framed as a search problem, where each node symbolizes a webpage. (a) ReACT: The agent chooses the best immediate actions without any proactive strategy. (b) Tree-search with reflections: the agent investigates different routes by actively navigating websites and allows for reversing direction (shown by dashed arrows). Both a) and b) approaches are Unknown-MDP-based. At each timestep, these agents’ observation space is constrained, which typically results in suboptimal or inefficient results. c) R2D2: Our proposed framework constructs the search space, leveraging stored state information from a replay buffer. By transforming the task into a Known MDP, R2D2 enhances its ability to navigate and interact with web interfaces.
Despite recent advancements in web agents’ capabilities, a persistent challenge remains: agents frequently fail to navigate effectively within intricate web environments as illustrated in Figure 1.
The fundamental challenges associated with previous methodologies are twofold. First, these approaches model web navigation as an Unknown Markov Decision Process (MDP), wherein the agent has limited visibility into the consequences of its actions, often leading to suboptimal performance outcomes. Second, prior methods engage in complex reasoning during the inference phase while observing a stream of experiences, and in their simplest forms, they discard incoming trajectories immediately after a single episode. This rapid forgetting of potentially valuable experiences impedes the agent’s capacity to leverage useful information for future decision-making.
@@
Inspired by cognitive studies showing that humans excel at complex tasks by iteratively refining strategies based on feedback Palenciano et al. (2021); Zenkri et al. (2024), as well as by approaches in robotics for structured exploration of unfamiliar spaces Thrun (2002), we propose the R2D2 (Remembering, Replaying, and Dynamic Decision Making) framework that enhances both navigation and task execution for web agents. Our method transforms the task from an Unknown-MDP into a Known MDP by introducing the Remember Paradigm. It leverages a structured replay buffer of the agent’s experiences that guides the agent to more promising avenues Blundell et al. (2016); Schaul et al. (2016); Parisotto and Salakhutdinov (2018); Savinov et al. (2018). At a high level, our approach enables the agent to record and recall previously visited pages—essentially constructing a dynamic map of the web environment—and then leverage this knowledge to refine its strategies. By converting the agent’s experience into a well-organized search space, we empower it to identify reliable navigation routes to target resources rather than re-deriving them from scratch during inference. This reduces the computational overhead at test time and helps avoid repetitive or unproductive exploration.
@@
Our proposed method diverges from traditional techniques by providing a more comprehensive and structured representation of the agent’s historical experiences. Instead of simply recalling past states Agashe et al. (2024) or relying on on-the-fly reasoning Koh et al. (2024); Zhou et al. (2024a), our approach organizes the agent’s accumulated experiences into a coherent and reusable resource that effectively guides future decisions. We evaluate our approach using the WebArena benchmark Zhou et al. (2024b), where it achieves substantial gains compared to baseline models, including approximately a 50% reduction in navigation errors and a threefold increase in overall task completion rates. Moreover, R2D2 outperforms state-of-the-art methods by 17%, thereby demonstrating a more robust and informed capability for executing complex web-based tasks.
@@ 2 Related Works
Enhancing Web Agents. Existing research has shown that language models, without intervention, struggle to express linguistic intent in formal instruction that can control an extra-linguistic environment, such as web site navigation Shi et al. (2017); Liu et al. (2018); Yao et al. (2022); Zhou et al. (2024b); Deng et al. (2023). This inadequacy stems primarily from the intrinsic challenges associated with perception, strategic planning, and task execution in the intricate web environments.
To mitigate these challenges, enhancements to web agents have been characterized as one of three principal strategies: (1) Perception Alignment: This strategy aims to augment agents’ capabilities in interpreting graphical user interface elements by integrating multimodal data from webpages, enhancing both textual and visual comprehension Gou et al. (2025); Liu et al. (2025). (2) Post-hoc Reflection: Studies indicate that enabling agents to engage in reflective practices post-interaction can facilitate learning from historical trajectories, thereby improving future task executions Shinn et al. (2023); Song et al. (2024); Pan et al. (2024); Wang et al. (2024a). (3) Online Search Algorithms: This involves the adoption of sophisticated search algorithms, including Monte Carlo Tree Search and other tree-based exploration methods, integrated with high-level planning driven by LLM-derived knowledge Koh et al. (2024); Meng et al. (2024); Zhang et al. (2025). Furthermore, Gu et al. 2024 discusses speculative planning that leverages simulations of world models.
Despite these enhancements, the performance of current web agents are constrained by the assumptions of Unknown MDP, where the potential outcomes of actions are not available. In contrast, this paper proposes a novel approach where we reconstruct the web environment’s structure based on an agent’s exploratory actions, thereby furnishing it with outcome information crucial for making informed and grounded decisions.
Continual Exploration of the Agentic Environment. Tasking agents to explore an unknown environment has been an active research direction Brohan et al. (2023). Recent studies have focused on how agents abstract experiences into actionable skills, a development that is becoming increasingly central to advancements in this field Wang et al. (2023, 2024c); Liu et al. (2024). Within the domain of web-based agents, these skills are often conceptualized as workflows. Sodhi et al. (2024) have introduced a novel framework that leverages human-engineered workflows to compose policies to tackle web tasks. Although this improves agent performance, manually crafting such workflows can be a tedious process.
@@
Figure 3: An overview of the R2D2 architecture, highlighting the Remember and Reflect Paradigms. The Remember paradigm constructs a structured replay buffer from previous observations, enabling the agent to use past episodic data through A* search for navigation. Meanwhile, the Reflect paradigm diagnoses errors and generates corrective insights, which are then stored in a reflective memory for future decision-making processes.
@@ 3 Method
In this section, we present our framework, which tackles complex web navigation tasks by integrating two paradigms: Remember and Reflect. The Remember paradigm constructs a structured replay buffer from past observations (Section 3.2), while the Reflect paradigm diagnoses and corrects errors in failed trajectories (Section 3.3). We then introduce mechanisms of the reflective memory (Section 3.4). Finally, we illustrate how these paradigms interact to improve the agent’s performance through retrieval and in-context learning demonstrations (Section 3.5).
@@ 3.1 Method Overview
Given a user’s task query $q$ and an initial observation $o_{0}$ from the environment, the agent must produce a sequence of actions to address $q$. We define an episode as the process where the agent starts from $o_{0}$ and executes a trajectory $t$. Let $t=\{a_{1},a_{2},\ldots,a_{H}\}$ be the trajectory of length $H$, where each $a_{h}$ is an action at step $h$. After each action $a_{h}$, the agent receives an observation $o_{h}$, thereby forming an observation sequence $O=\{o_{1},o_{2},\ldots,o_{H}\}$. Consider $N$ distinct user queries $\{q_{1},q_{2},\ldots,q_{N}\}$, each associated with its own episode and observation sequence $O_{i}=\{o_{i,1},o_{i,2},\ldots,o_{i,H}\}$. Across all $N$ user queries, R2D2 forms the union of these observations:$\mathcal{O}_{\text{all}}=\bigcup_{i=1}^{N}O_{i}.$
R2D2 addresses errors in trajectories by categorizing them into two distinct types: (1) Navigation failure. We define $O^{*}$ as the key observations essential for successfully addressing the query $q$. After performing the trajectory $t$, the observation sequence may not contain all the key observations, which leads to agents’ navigation failure in the environment, formally $O\cap O^{*}\neq O^{*}$. The agent fails because of incomplete information or tools to address the user query. (2) Execution failure. After performing $t$, $O\cap O^{*}=O^{*}$. In other words, navigation was successful, since the proper path was followed, but the agent still failed to address the user query.
The Remember paradigm aims to build a replay buffer from $\mathcal{O}_{all}$, allowing the agent to store and revisit past observations and experiences Blundell et al. (2016); Schaul et al. (2016); Parisotto and Salakhutdinov (2018); Savinov et al. (2018). The Reflection paradigm then corrects trajectories that failed due to execution issues, providing explicit rationales for these execution failures. The successful and corrected trajectories and their rationales are stored in a reflective memory for the agent’s future reference. Finally, the Retriever leverages this reflective memory by selecting relevant corrected trajectories as in-context demonstrations, thereby continually improving the agent’s performance.
@@ 3.2 Remember Paradigm
R2D2 first determines [^2] Please refer to Figure 7 in Appendices for the detailed prompt. the error type of a failed trajectory $t$. For all navigation failures, R2D2 performs search within the replay buffer to correct their navigation behaviors.
Building the Replay Buffer. To effectively represent the web environment and enable its systematic reconstruction from the observation sequence $\mathcal{O}_{\text{all}}$, R2D2 structures the replay buffer as a directed graph $G=(O,E)$. Here, the vertex set $O$ includes the root node $o_{0}$, which corresponds to the homepage observation, and each subsequent vertex represents another webpage observation $o_{i}$. The edge set $E$ consists of tuples $((o_{i},o_{j}),a)$ where each edge corresponds to an action $a$ that transitions the agent from one observation $o_{i}$ to another observation $o_{j}$. Due to the noisy and dynamic nature of web pages, R2D2 stores the differences between consecutive observations at each vertex rather than the full webpage state.
$\mathbf{A}^{*}$ Search within the Buffer. Our search algorithm employs a best-first search ($A^{*}$) strategy Hart et al. (1968); Meng et al. (2024) to navigate and evaluate web environments effectively. Instead of expanding nodes level-by-level, R2D2 incorporates a heuristic that guides the search toward potentially relevant and promising webpages more efficiently Bonet and Geffner (1999); Guez et al. (2018); Moldovan and Abbeel (2012). This heuristic, provided by the LLM, estimates the relevance and utility of exploring a particular webpage node, thereby reducing unnecessary expansions and focusing on paths that are more likely to yield correct information.
@@
0:  Optimal solution trajectory $P^{*}$ for $q$
@@ 3.3 Reflect Paradigm
We discussed details of how we address navigation failures in Section 3.2. We now discuss using reflection techniques to address execution errors.
@@ 3.4 Reflective Memory Mechanism
We introduce the reflective memory mechanism that stores corrected and truncated trajectories for future retrieval. The reflective memory is structured as a key-value store:
@@ 3.5 Paradigm Coordination and Inference
Exploration Phase. Using a ReACT agent Yao et al. , R2D2 processes user queries and collects observational data to build $\mathcal{O}_{\text{all}}$. Trajectories are classified and corrected via Remember and Reflect paradigms, then stored in the memory.
@@ 4.1 Experimental Setup
ReACTYao et al. 13.1%
Tree-Search Koh et al. (2024) 17% 11% 28% 26% 13% 19.0%
@@
Benchmark. We use the WebArena benchmark Zhou et al. (2024b). This benchmark comprises diverse web interaction scenarios, ranging from web shopping to customer relationship management system (CMS). The dataset consists of 812 user queries with annotated ground truth trajectories. The Webarena benchmark further provides a set of validators to programmatically validate the functional correctness of each task.
@@
Baselines. We compare our framework against several representative agent frameworks: (1) ReACT Yao et al. : a widely-used framework, which takes an observation of the environment as input, performs Chain-of-Thought reasoning, and then generates the next action. (2) Tree-Search Koh et al. (2024): an inference-time tree-search strategy to perform best-first tree-search in web environments. It enables agents to revert to the most recently validated state upon encountering a failed trajectory. (3) LATS Zhou et al. (2024a): a method based on Monte Carlo tree search that employs LLMs as agents, value functions, and optimizers for decision-making. (4) Anticipatory Reflection Wang et al. (2024a): a method that explicitly considers potential failures before action, alignment and backtracking after actions to maintain task objectives. (5) AutoEval Pan et al. (2024): methods that boost agent performance using domain-general automatic evaluators. (6) BrowserGym Chezelles et al. (2024): a framework that incorporates additional actions and observation tools for agents to interact with the environment. [^5] There are other methods that use different setups, such as SteP Sodhi et al. (2024), that employs human-engineered workflows, and AWM Wang et al. (2024b), which uses BrowserGym framework and customizes a larger action space than standard WebArena. For the sake of direct comparison, these frameworks that depend on additional human efforts are not taken into comparison here.